A couple of textbook questions I'm having a tough time answering:

1.) Suppose​ that:

r​ = required reserve ratio​ = 0.10

c ​= ​{C/D}​ = currency ratio​ = 0.45

e​ = ​{ER/D}​ = excess reserve ratio​ = 0.03

MB​ = the monetary base​ = $3000 billion

Given that the formula for the money multiplier is (1+c/r+e+c) find the value for M​, the money supply.

The money supply is $____ billion.​ (Round your response to the nearest whole​ number.)

Use the money multiplier to find the new value for the money supply if open market operations increase the monetary base by $100 billion.

The money supply is now $____ billion. ​(Round your response to the nearest whole​ number.)

2.) The purpose of the commitment by the Fed to keep the federal funds rate at zero for a long period of time is to:

a.) Lower the short term interest rates

b.) Lower the long term interest rates

c.) Increase the short term interest rates

d.) Increase the long term interest rate

3.) Disadvantages of the Fed's "just do it" approach include:

a.) a lack of transparency, which creates doubt about the future course of inflation and output and makes it hard to hold the Fed accountable

b.) the strong dependence on the preferences, skills, and the trustworthiness of the individuals in charge of the central bank

c.) low accountability that my make the Fed more susceptible to the time-inconsistency problem

d.) Only A and B are correct

e.) All of the above are correct

(note: I'm fairly certain D is the correct answer here, but just wanted some verification)

Thank you!
Answer #1


Question 1

r = 0.10

c = 0.45

e = 0.03

Calculate the money multiplier -

m = (1+c)/(r+e+c)

m = (1+0.45)/(0.10+0.03+0.45)

m = 1.45/0.58

m = 2.5

The money multiplier is 2.5

Calculate the money supply -

Money Supply = Money Multiplier * Monetary Base

Money Supply = 2.5 * $3,000 billion = $7,500 billion

The money supply is $7,500 billion.

Now, monetary base increases by $100 billion.

So, monetary base will now be ($3,000 billion + $100 billion) $3,100 billion.

Calculate the money supply -

Money supply = Money Multiplier * Monetary Base

Money supply = 2.5 * $3,100 billion = $7,750 billion

The money supply is now $7,750 billion.